23-t< or = 2(t-9)-3(t+2)
solve for T

Answer by stanbon(75887)

23-t< or = 2(t-9)-3(t+2)
solve for T
-------------------------
23 - t <= 2(t-9)-3(t+2)
23 - t <= 2t-18-3t-6
23 - t <= -t-24
-----------
Adding t to both sides you get
23<=-24
This is wrong.
---
The inequality is inconsistent.
It has no solution.
=============================
Cheers,
Stan H.